OpenStudy (anonymous):

Four sheep are traveling the 250 m distance from the pasture back to the barn. After 45 seconds, their distances from the barn are as follows: Dolly - 170 m, Fluffy - 115 m, Ellie - 90 m, Ma - 200 m. According to average velocity, how would the sheep rank from slowest to fastest?

OpenStudy (anonymous):

Traveled Average velocity Dolly (250 - 170 = 80) 80*(4/3) = 320/3 mile/hr Fluffy (250 - 115 = 135) 135*(4/3) = 540/3 Ellie (250 - 90 = 160) 160*(4/3) = 640/3 Ma (250 - 200 = 50) 50*(4/3) = 200/3 Slowest to fastest: Ma - Dolly - Fluffy - Ellie